# Sell It Before You Make It: Revolutionizing E-Commerce with Personalized AI-Generated Items

### v1

@@ Abstract.
E-commerce has revolutionized retail, yet its traditional workflows remain inefficient, with significant time and resource costs tied to product design and manufacturing inventory. This paper introduces a novel system deployed at Alibaba that leverages AI-generated items (AIGI) to address these challenges with personalized text-to-image generation for e-commercial product design. AIGI enables an innovative business mode called “sell it before you make it”, where merchants can design fashion items and generate photorealistic images with digital models based on textual descriptions. Only when the items have received a certain number of orders, do the merchants start to produce them, which largely reduces reliance on physical prototypes and thus accelerates time to market. For such a promising application, we identify the underlying key scientific challenge, i.e., capturing the users’ group-level personalized preferences towards multiple generated candidate images. To this end, we propose a Personalized Group-Level Preference Alignment Framework for Diffusion Models (i.e., PerFusion). We first design PerFusion Reward Model for user preference estimation with a feature-crossing-based personalized plug-in. Then we develop PerFusion with a personalized adaptive network to model diverse preferences across users, and meanwhile derive the group-level preference optimization objective to capture the comparative behaviors among multiple candidates. Both offline and online experiments demonstrate the effectiveness of our proposed algorithm. The AI-generated items have achieved over 13% relative improvements for both click-through rate and conversion rate compared to their human-designed counterparts, validating the revolutionary potential of AI-generated items for e-commercial platforms.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
Figure 1. The illustration of (a) traditional procedure and (b) innovative procedure with AI-generated items (AIGI) for e-commercial platforms.
@@ 1. Introduction
This paper introduces a system deployed at Alibaba that presents a pioneering attempt at realizing the transformative potential of generative AI in the e-commerce industry. By integrating cutting-edge generative text-to-image models, we automate the fashion product design process and validate the innovative “sell it before you make it” business mode in a real-world e-commercial environment. As shown in Figure 2, the core workflow of our system begins with the user (i.e., the merchant or designer) providing a textual prompt about the desired product design. Once the prompt is submitted, a diffusion model generates a diverse set of candidate images. The user then evaluates these candidates, selecting $K$ preferred images from the $N$ generated options. If none of the images meet the user’s expectations, he/she can refine the text prompt and initiate another round of generation. This iterative process empowers users to progressively refine the item designs to align with their specific tastes.
@@
Firstly, group-level comparative preference means that user preferences are expressed by comparing a group of images rather than through individual or pairwise evaluations. As shown in Figure 2, each user should assess the entire group of candidate images holistically to make the decision. However, existing preference alignment algorithms for diffusion models either solely rely on positive images without adequately considering negatives (Ho et al., 2020), or focus on pairwise optimization based on one pair of positive and negative images at a time (Wallace et al., 2024; Croitoru et al., 2024; Liang et al., 2024). These methods are not well suited for our application and would lose the global view of the group-level context, thereby leading to suboptimal performance.
Secondly, personalized preference across users indicates that user preferences exhibit significant variability, even when the same textual prompt and image set are provided. This indicates that textual prompts alone are insufficient to fully encapsulate the implicit visual preferences of individual users. There are preliminary works exploring user-level preference optimization by incorporating either textual reconstruction (Yang et al., 2024; Czapp et al., 2024) or personalized classifier-free guidance with delicately designed networks (Xu et al., 2024b, d). However, they are generally built to re-paint or re-design the reference image based on certain existing contents, e.g., product background outpainting (Shilova et al., 2023; Yang et al., 2024), poster re-painting (Xu et al., 2024d; Shen et al., 2024), and fashion collocation (Xu et al., 2024b). When faced with the task of generating realistic images from scratch, capturing subtle variations in users’ visual preferences still remains a significant challenge.
To this end, in this paper, we propose a Personalized Group-Level Preference Alignment Framework for Diffusion Models (i.e., PerFusion). We first design the PerFusionRM (short for PerFusion Reward Model) for personalized user preference estimation based on the CLIP model (Radford et al., 2021). We propose a feature-crossing-based personalized plug-in to extract the user preference representations and inject them into each Transformer layer of CLIP. The design of PerFusionRM serves as the preliminary step to verify the fact that the personalized knowledge can truly affect the user’s decision for text-to-image generation. Moreover, PerFusionRM is also adopted as the reward model to evaluate the performance of personalized text-to-image generation. Then, we propose the PerFusion framework based on the stable diffusion (Rombach et al., 2022) architecture for image generation. Specifically, we design a personalized adaptive network to guide the diffusion model with additional user preference conditions, capturing the diverse preferences across different individual users. Moreover, a group-level preference optimization objective is derived to model the comparative behavior patterns of users towards multiple candidate images.
@@
  • We propose the PerFusion framework to capture the users’ group-level personalized preferences. We first design PerFusion Reward Model for user preference estimation with a feature-crossing-based personalized plug-in. Then we develop PerFusion based on stable diffusion (Rombach et al., 2022) with a personalized adaptive network to model diverse preferences across users, and meanwhile derive the group-level preference optimization objective to capture the comparative behaviors among multiple candidate images.
  • Both offline and online experiments show the effectiveness of our proposed algorithm. The AI-generated items have achieved over 13% relative improvements for both click-through rate and conversion rate compared with their human-designed counterparts. This validates the revolutionary potential of the “sell it before you make it” business mode based on AI-generated items (AIGI).
@@ 2. Related Works
2.1. Image Generation for E-Commerce
Image generation in e-commerce aims to customize item displays based on user preferences and maintain visual consistency to enhance the user experience. Generative models offer automated and personalized content creation of product images, thumbnails, background visuals, and more (Yu et al., 2019; Lin et al., 2024; Wei and Fang, 2025; Hu et al., 2024; Kim et al., 2024; Vashishtha et al., 2024; Wang et al., 2023). These advancements have expanded the application of image generation in e-commerce to areas such as fashion generation and advertisement creative generation. Fashion generation (Kang et al., 2017; Shih et al., 2018; Yang et al., 2018; Xu et al., 2024c; Rombach et al., 2022) involves generating fashion-related contents for personalized fashion recommendations and outfit collocation. Advertisement creative generation (Kanungo et al., 2021; Mita et al., 2023; Shilova et al., 2023; Yang et al., 2024; Czapp et al., 2024; Vashishtha et al., 2024) focuses on producing engaging advertising content, like promotional fonts, backgrounds, and banners tailored for different audiences.
However, current image generation in e-commerce relies on existing items and physical prototypes. In contrast, we are the first to propose and deploy a real-world industrial image generation service for AI-generated items (AIGI) in e-commerce. In this paper, we not only address the scientific problem of capturing users’ group-level personalized preferences for image generation, but also enable the innovative “sell it before you make it” business mode.
2.2. Preference Alignment for Diffusion Models
The preference alignment for diffusion models can be categorized into (1) general preference alignment and (2) personalized preference alignment. General preference alignment (Karthik et al., 2024; Wallace et al., 2024; Liang et al., 2024; Ho et al., 2020) optimizes the text-image consistency and common aesthetic appeal, helping models adhere to visual principles like realism, coherence, and alignment with textual prompts. While effective, these approaches only concentrate on the general human preference annotated by crowdsourcing or general reward models, and thereby fail to capture the nuanced preferences of different individual users. Hence, personalized preference alignment (Xu et al., 2024c; Mukande et al., 2024; Czapp et al., 2024; Xu et al., 2024d; Wei and Fang, 2025) emerges as a promising solution, especially for online services like e-commerce. It adapts generated outputs to individual aesthetic tastes, which are often implicit and difficult to express via text prompts.
Compared with existing works, the core difference and contribution of this paper are in two folds. (1) We are the first to explicitly model the group-level personalized preference for diffusion models, which jointly considers the group-level comparative pattern and personalized preference across users. (2) We have deployed the image generation services in Alibaba, demonstrating the potential of AI-generated items (AIGI) without existing products or physical prototypes for e-commercial platforms.
@@ 3.1. Problem Formulation
We aim to model users’ group-level personalized preferences during the image generation process for our e-commercial platforms. Specifically, we consider a user, typically an e-commerce merchant in our application, who has a set of $F$ features denoted as =$\mathcal{F}=\{f_{i},f_{2},...,f_{F}\}$, e.g., product style and price level. At each round of image generation, the user inputs the text description $txt$ and is provided with a group of generated candidate images $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$. Then, the user gives the feedback $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$, where $y_{i}\in\{1,0\}$ indicates whether the user selects (i.e., prefers) the image $x_{i}$ or not. Note that each group of images is represented to the user at the same time, thereby showcasing the group-level comparative preference of the user. Without loss of generality, we assume that different image groups have the same group size of $N$. Given this setup, we decompose the goal of modeling users’ group-level personalized preferences into two folds:
  • User Preference Estimation: Given a textual instruction $txt$ and a group of candidate images $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, we aim to accurately predict the user’s preference for each image within the group, i.e., $\{\hat{y}_{i}\}_{i=1}^{N}$. We should not only measure the consistency between the textual prompt and generated images, but also capture the user’s visual preference among the image group.
  • Personalized Image Generation: Given a textual instruction $txt$, we aim to generate an image that best aligns with the user’s preferences, i.e., $\hat{x}$. The generated image should not only reflect the given description but also meet the implicit visual preferences inferred from the user’s profile features.
@@ 3.2. Diffusion Models
Since our proposed method and application scenarios are built based on stable diffusion (Rombach et al., 2022) (SD), we provide a brief overview of its underlying principles. Stable diffusion is a latent diffusion model (LDM) that generates high-fidelity images by iteratively refining Gaussian noise through a denoising process. Unlike conventional diffusion models operating directly in pixel space (Ho et al., 2020), SD performs diffusion in a latent space, necessitating an autoencoder comprising an encoder and a decoder.
We denote the encoder as $\mathcal{E}(\cdot)$, which maps an image $x$ to its latent representation $z=\mathcal{E}(x)$. Conversely, the decoder, denoted as $\mathcal{D}(\cdot)$, reconstructs the image from its latent representation. Given an input text prompt $txt$, the text encoder $\mathcal{C}(\cdot)$ encodes it into a text embedding $c=\mathcal{C}(txt)$, which serves as the conditioning input for stable diffusion. During training, for a given diffused latent representation $z_{t}$ at timestep $t$ with condition $c$, the denoising network $\epsilon_{\theta}(\cdot)$ is trained to predict the noise $\epsilon$. This is achieved by minimizing the following objective function:
@@
where $\epsilon_{\theta}(\cdot)$ can be typically implemented using the U-Net (Ronneberger et al., 2015) or Transformer (Vaswani, 2017) architecture.
@@ 4. Methodology
In this section, we introduce our Personalized Group-level Preference Alignment Framework for Diffusion Models (i.e., PerFusion).
4.1. Overview
@@
We first design the PerFusionRM (short for PerFusion Reward Model) for personalized user preference estimation based on the CLIP model (Radford et al., 2021; Cherti et al., 2023). Specifically, we propose a feature-crossing network to extract the user preference knowledge from the profile features, and then inject the knowledge into each Transformer layer of the CLIP model. The feature-crossing network serves as a personalized plug-in, which can retain the general preference estimation of the original CLIP model and leverage the personalized information for adaptation.
@@
Then, we introduce the PerFusion framework to diffusion models for personalized image generation. To be specific, we adopt the similar feature-crossing network plug-in from the PerFusionRM to extract the user’s preference knowledge, which provides personalized conditions for Stable Model through a ControlNet (Zhang et al., 2023) manner. We also derive the group-level preference optimization objective to capture the comparative behavior patterns.
@@ 4.2. User Preference Estimation
Specifically, given the user features =$\mathcal{F}=\{f_{1},\dots,f_{F}\}$, we first apply the embedding layer to map them to low-dimensional dense embedding vectors $\{v1,\dots,v_{F}\}$. Then, following previous research on feature interaction (Zhang et al., 2021; Wang et al., 2021, 2017), we apply a multi-layer feature crossing network with residual connections to extract the co-occurrence and co-relation patterns among the features:
@@
(3) =${}_{i=1}^{T}=\text{MHA}\left([h_{l,i}]_{i=1}^{T}\right),\;\;[h_{l+1,i}]_{i=1}^{T}=\text{FFN}\left([h_{l,i}^{\prime}]_{i=1}^{T}\right),$
@@
where $L$ is the number of Transformer layers for the CLIP tower, and the adaptive networks are implemented as two-layer perceptrons with GeLU activations (Hendrycks and Gimpel, 2016). Then, we insert the adaptive knowledge $\{u_{\text{attn}}^{l},u_{\text{ffn}}^{l}\}$ into the Transformer layer with element-wise addition:
@@
(6) $s(txt,x)=\text{Cosine\_Similarity}\left(v_{\text{text}}^{u},v_{\text{image}}^{u}\right).$
Each training sample can be formulated as $\{txt,\mathcal{X},\mathcal{F},\mathcal{Y}\}$, where $txt$ is text description, $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ is the group of candidate images, $\mathcal{F}$ is the user feature set, and $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$ is the set of corresponding binary labels. We can compute the ideal preference distribution as:
(7) $p_{i}=\frac{y_{i}}{\sum_{j=1}^{N}y_{j}},\;\,i=1,\dots,N.$
For instance, if the user prefers the first two images and rejects all the rest, then the ideal preference distribution is $[0.5,0.5,0,\dots]$. Similarly, the predicted distribution can be calculated as:
@@
Note that we freeze the original CLIP backbone and solely finetune the parameters of personalized plug-in modules, which helps retain the basic knowledge for estimating text-image pairs.

### v2

@@ Abstract.
E-commerce has revolutionized retail, yet its traditional workflows remain inefficient, with significant resource costs tied to product design and inventory. This paper introduces a novel system deployed at Alibaba that uses AI-generated items (AIGI) to address these challenges with personalized text-to-image generation for e-commerce product design. AIGI enables an innovative business mode called “sell it before you make it”, where merchants can design fashion items and generate photorealistic images with digital models based on textual descriptions. Only when the items have received a certain number of orders, do the merchants start to produce them, which largely reduces reliance on physical prototypes and thus accelerates time to market. For such a promising application, we identify the underlying key scientific challenge, i.e., capturing users’ group-level personalized preferences towards multiple generated images. To this end, we propose a Personalized Group-Level Preference Alignment Framework for Diffusion Models (PerFusion). We first design PerFusion Reward Model for user preference estimation with a feature-crossing-based personalized plug-in. Then we develop PerFusion with a personalized adaptive network to model diverse preferences across users, and meanwhile derive the group-level preference optimization objective to model comparative behaviors among multiple images. Both offline and online experiments demonstrate the effectiveness of our proposed algorithm. The AI-generated items achieve over 13% relative improvements for both click-through rate and conversion rate, as well as 7.9% decrease in return rate, compared to their human-designed counterparts, validating the transformative potential of AIGI for e-commerce platforms.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783918isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Personalizationccs: Information systems Multimedia content creationccs: Applied computing Electronic commerce
Figure 1. The illustration of (a) traditional procedure and (b) innovative procedure with AI-generated items (AIGI) for e-commerce platforms.
@@ 1. Introduction
This paper introduces a system deployed at Alibaba that presents a pioneering attempt at realizing the transformative potential of generative AI in the e-commerce industry. By integrating cutting-edge generative text-to-image models, we automate the fashion product design process and validate the innovative “sell it before you make it” business mode in a real-world e-commerce environment. As shown in Figure 2, the core workflow of our system begins with the user (i.e., the merchant or designer) providing a textual prompt about the desired product design. Once the prompt is submitted, a diffusion model generates a diverse set of candidate images. The user then evaluates these candidates, selecting $K$ preferred images from the $N$ generated options. If none of the images meet the user’s expectations, the user can refine the text prompt and initiate another round of generation. This iterative process empowers users to progressively refine the item designs to align with their specific tastes. It is crucial to note that our system is a merchant-facing tool; the end customers on the e-commerce platform are generally unaware of whether a product is AI-generated, ensuring a seamless shopping experience.
@@
Firstly, group-level comparative preference means that user preferences are expressed by comparing a group of images rather than through individual or pairwise evaluations. As shown in Figure 2, each user should assess the entire group of candidate images holistically to make the decision. However, existing preference alignment algorithms for diffusion models either solely rely on positive images without adequately considering negatives (ho2020denoising), or focus on pairwise optimization based on one pair of positive and negative images at a time (wallace2024diffusion; croitoru2024curriculum; liang2024step). These methods are not well suited for our application and would lose the global view of the group-level context, thereby leading to suboptimal performance.
Secondly, personalized preference across users indicates that user preferences exhibit significant variability, even when the same textual prompt and image set are provided. This indicates that textual prompts alone are insufficient to fully encapsulate the implicit visual preferences of individual users. There are preliminary works exploring user-level preference optimization by incorporating either textual reconstruction (yang2024new; czapp2024dynamic) or personalized classifier-free guidance with delicately designed networks (xu2024difashion; xu2024personalized). However, they are generally built to re-paint or re-design the reference image based on certain existing contents, e.g., product background outpainting (shilova2023adbooster; yang2024new), poster re-painting (xu2024personalized; shen2024pmg), and fashion collocation (xu2024difashion). When faced with the task of generating realistic images from scratch, capturing subtle variations in users’ visual preferences still remains a significant challenge.
@@
To this end, in this paper, we propose a Personalized Group-Level Preference Alignment Framework for Diffusion Models (i.e., PerFusion). We first design the PerFusionRM (short for PerFusion Reward Model) for personalized user preference estimation based on the CLIP model (radford2021learning). We propose a feature-crossing-based personalized plug-in to extract the user preference representations and inject them into each Transformer layer of CLIP. The design of PerFusionRM serves as the preliminary step to verify the fact that the personalized knowledge can truly affect the user’s decision for text-to-image generation. Moreover, PerFusionRM is also adopted as the reward model to evaluate the performance of personalized text-to-image generation. Then, we propose the PerFusion framework based on the stable diffusion (rombach2022high) architecture for image generation. Specifically, we design a personalized adaptive network to guide the diffusion model with additional user preference conditions, capturing the diverse preferences across different individual users. Moreover, a group-level preference optimization objective is derived to model the comparative behavior patterns of users towards multiple candidate images.
@@
  • We propose the PerFusion framework to capture the users’ group-level personalized preferences. We first design PerFusion Reward Model for user preference estimation with a feature-crossing-based personalized plug-in. Then we develop PerFusion based on stable diffusion (rombach2022high) with a personalized adaptive network to model diverse preferences across users, and meanwhile derive the group-level preference optimization objective to capture the comparative behaviors among multiple candidate images.
  • Both offline and online experiments show the effectiveness of our proposed algorithm. This validates the revolutionary potential of the “sell it before you make it” business mode based on AI-generated items (AIGI).
@@ 2. Related Works
Image Generation for E-Commerce. Image generation in e-commerce aims to customize item displays based on user preferences and maintain visual consistency to enhance the user experience (lin2025diffusion). Generative models offer automated and personalized content creation of product images, thumbnails, background visuals, and more (yu2019personalized; lin2024survey; wei2025diffusion; hu2024fashionr2r; kim2024stableviton; vashishtha2024chaining; wang2023generative). These advancements have expanded the application of image generation in e-commerce to areas such as fashion generation and advertisement creative generation. Fashion generation (kang2017visually; shih2018compatibility; yang2018recommendation; xu2024diffusion; rombach2022high) involves generating fashion-related contents for personalized fashion recommendations and outfit collocation. Advertisement creative generation (kanungo2021ad; mita2023camera; shilova2023adbooster; yang2024new; czapp2024dynamic; vashishtha2024chaining) focuses on producing engaging advertising content, like promotional fonts, backgrounds, and banners tailored for different audiences.
Preference Alignment for Diffusion Models. The preference alignment for diffusion models can be categorized into (1) general preference alignment and (2) personalized preference alignment. General preference alignment (karthik2024scalable; wallace2024diffusion; liang2024step; ho2020denoising) optimizes the text-image consistency and common aesthetic appeal, helping models adhere to visual principles like realism, coherence, and alignment with textual prompts. While effective, these approaches only concentrate on the general human preference annotated by crowdsourcing or general reward models, and thereby fail to capture the nuanced preferences of different individual users. Hence, personalized preference alignment (xu2024diffusion; mukande2024mmcrec; czapp2024dynamic; xu2024personalized; wei2025diffusion) emerges as a promising solution, especially for online services like e-commerce. It adapts generated outputs to individual aesthetic tastes, which are often implicit and difficult to express via text prompts.
@@ 3.1. Problem Formulation
We aim to model users’ group-level personalized preferences during the image generation process for our e-commerce platforms. Specifically, we consider a user, typically an e-commerce merchant in our application, who has a set of $F$ features denoted as $\mathcal{F}=\{f_{i}\}_{i=1}^{F}$, e.g., shop style and price level. At each round of image generation, the user inputs the text description txttxt and is provided with a group of generated candidate images $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$. Then, the user gives the feedback $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$, where $y_{i}\in\{1,0\}$ indicates whether the user selects (i.e., prefers) the image $x_{i}$ or not. Note that each group of images is represented to the user at the same time, thereby showcasing the group-level comparative preference of the user. We assume that different image groups have the same group size of $N$. Given this setup, we decompose the goal of modeling users’ group-level personalized preferences into two folds:
  • User Preference Estimation: Given a textual instruction txttxt and a group of candidate images $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, we aim to accurately predict the user’s preference for each image within the group, i.e., $\{\hat{y}_{i}\}_{i=1}^{N}$. We should not only measure the consistency between the textual prompt and generated images, but also capture the user’s visual preference among the image group.
  • Personalized Image Generation: Given a textual instruction txttxt, we aim to generate an image that best aligns with the user’s preferences, i.e., $\hat{x}$. The generated image should not only reflect the given description but also meet the implicit visual preferences inferred from the user’s profile features.
@@ 3.2. Diffusion Models
Our method is built based on stable diffusion (rombach2022high) (SD). SD is a latent diffusion model (LDM) that generates high-fidelity images by iteratively refining Gaussian noise through a denoising process. Unlike conventional diffusion models operating directly in pixel space (ho2020denoising), SD performs diffusion in a latent space, necessitating an autoencoder comprising an encoder and a decoder.
We denote the encoder as $\mathcal{E}(\cdot)$, which maps an image $x$ to its latent representation $z=\mathcal{E}(x)$. Conversely, the decoder, denoted as $\mathcal{D}(\cdot)$, reconstructs the image from its latent representation. Given an input text prompt txttxt, the text encoder $\mathcal{C}(\cdot)$ encodes it into a text embedding $c=\mathcal{C}(txt)$, which serves as the conditioning input for stable diffusion. During training, for a given diffused latent representation $z_{t}$ at timestep $t$ with condition $c$, the denoising network $\epsilon_{\theta}(\cdot)$ is trained to predict the noise $\epsilon$. This is achieved by minimizing the following objective function:
@@
where $\epsilon_{\theta}(\cdot)$ can be typically implemented using the U-Net (ronneberger2015u) or Transformer (vaswani2017attention) architecture.
@@ 4. Methodology
4.1. Overview of PerFusion
@@
We first design the PerFusionRM (short for PerFusion Reward Model) for personalized user preference estimation based on the CLIP model (radford2021learning; cherti2023reproducible). Specifically, we propose a feature-crossing network to extract the user preference knowledge from the profile features, and then inject the knowledge into each Transformer layer of the CLIP model. The feature-crossing network serves as a personalized plug-in, which can retain the general preference estimation of the original CLIP model and leverage the personalized information for adaptation.
@@
Then, we introduce the PerFusion framework to diffusion models for personalized image generation. To be specific, we adopt the similar feature-crossing network plug-in from the PerFusionRM to extract the user’s preference knowledge, which provides personalized conditions for Stable Model through a ControlNet (zhang2023adding) manner. We also derive the group-level preference optimization objective to capture the comparative behavior patterns.
@@ 4.2. User Preference Estimation
Specifically, given the user features $\mathcal{F}=\{f_{1},\dots,f_{F}\}$, we first apply the embedding layer to map them to low-dimensional dense embedding vectors $\{v1,\dots,v_{F}\}$. Then, following previous research on feature interaction (wang2021dcn; wang2017deep), we apply a multi-layer feature crossing network with residual connections to extract the co-occurrence and co-relation patterns among the features:
@@
(3) ${}_{i=1}^{T}=\text{MHA}\left([h_{l,i}]_{i=1}^{T}\right),\;\;[h_{l+1,i}]_{i=1}^{T}=\text{FFN}\left([h_{l,i}^{\prime}]_{i=1}^{T}\right),$
@@
where $L$ is the number of Transformer layers for the CLIP tower, and the adaptive networks are implemented as two-layer perceptrons with GeLU activations (hendrycks2016gaussian). Then, we insert the adaptive knowledge $\{u_{\text{attn}}^{l},u_{\text{ffn}}^{l}\}$ into the Transformer layer with element-wise addition:
@@
(6) $s(txt,x)=\text{Cosine\_Similarity}(v_{\text{text}}^{u},v_{\text{image}}^{u}).$
Each training sample can be formulated as $\{txt,\mathcal{X},\mathcal{F},\mathcal{Y}\}$, where txttxt is text description, $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ is the group of candidate images, $\mathcal{F}$ is the user feature set, and $\mathcal{Y}=\{y_{i}\}_{i=1}^{N}$ is the set of corresponding binary labels. We can compute the ideal preference distribution as:
(7) $p_{i}=y_{i}\;/\;\sum\nolimits_{j=1}^{N}y_{j},\;\,i=1,\dots,N.$
For instance, if the user prefers the first two images and rejects all the rest, then the ideal preference distribution is $0.5,0.5,0,\dots]$. Similarly, the predicted distribution can be calculated as:
@@
This objective enables the injection of rich, structured merchant-side information into the reward model, enabling fine-grained preference learning. Note that we freeze the original CLIP backbone and solely finetune the parameters of personalized plug-in modules, which helps retain the basic knowledge for estimating text-image pairs.